# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ Abstract
Symmetry under a particular class of non-strictly canonical transformations may be used to identify, and subsequently excise degrees of freedom which do not contribute to the closure of the algebra of dynamical observables. In this article, we present a mathematical framework which extends this symmetry reduction procedure to classical field theories described within the covariant De-Donder Weyl formalism. Both the Lagrangian and Hamiltonian formalisms are discussed; in the former case, the Lagrangian is introduced as a bundle morphism $\mathcal{L}:J^{1}E\rightarrow\wedge^{m}M$, whilst the Hamiltonian description takes as its multiphase space the restricted multimomentum bundle $J^{1}E^{*}$. The contact reduction process is implemented in a number of particular cases, and it is found that the geometry of the reduced space admits an elegant physical interpretation.
@@ I Introduction
When attempting to formulate a symplectic description of classical field theories, there exist two distinct approaches, both of which have proved highly fruitful in their respective areas of application. On the one hand, the spacetime-covariant formalism, described within the framework of multisymplectic geometry, treats events at a single spacetime point, considering all coordinates (including time) equally [1, 2]; the canonical (or $3+1$) construction, by contrast, aims to study the evolution of a dynamical system from the perspective of the initial-value problem, defined on some suitably-chosen Cauchy surface [3, 4]. In this article, we shall almost exclusively be concerned with the covariant formalism, which in the context of the Hamiltonian description, we refer to as the Hamilton De-Donder Weyl (HDW) formulation [5, 6].
The use of multisymplectic geometry to describe classical field theories presents a number of advantages over its canonical counterpart; in particular, the instantaneous Hamiltonian formalism, almost ubiquitously employed as a precursor to quantum field theory, takes as its canonical variables field functions, evaluated at some fixed instant of time [7]. Consequently, the resulting phase space, upon which we are obliged to work, is infinite-dimensional [8]. Additionally, since the HDW formalism makes no distinction amongst spacetime coordinates when constructing conjugate momenta [9], it is, by its very nature, manifestly covariant [10]. These two features (finite-dimensionality and manifest covariance) have led to significantly hightened research interest within the field of multisymplectic geometry in recent years.
In spite of this, the geometrical formulation is certainly not without its challenges; it is still not fully understood, for example, how the Poisson bracket should be constructed, which, naturally, is an essential prerequisite for most potential geometric quantisation scheme [11, 12, 13]. Some progress has been made in this field [14, 15], and a multisymplectic bracket is well-defined between semi-basic forms on the covariant phase space [16]; however, a fully rigorous construction, applicable to arbitrary forms remains, for the most part, an open question. Additionally, the equal treatment of all spacetime coordinates - which confers the attractive quality of manifest covariance - leads to a discrepancy in the number of field and momentum degrees of freedom: to each field variable, there correspond as many conjugate momenta as dimensions of the underlying spacetime manifold.
Our most precise and accurate description of nature is currently formulated in the language of field theory, both at a classical and quantum level [17]. Indeed, the advent of a field-theoretic viewpoint has arguably been one of the most significant shifts in perspective of the last few centuries. It has been argued by philosophers such as Ismael and van Fraassen that, when seeking a theoretical description of the natural world, one should first construct a framework which amply describes the phenomena of interest, before progressively reducing its superfluous structure, until a minimally-sufficient yet consistent theory remains [18, 19, 20]. It is therefore of interest to examine the structural elements of our field-theoretic models which may be deemed dispensable, and further, to explore any novel insights which are made manifest, as a result of adopting this minimalistic viewpoint. To this end, it is necessary to formulate a precise and rigorous prescription for the identification and subsequent elimination of any structure which is not truly fundamental to our theory.
Our characterisation of mathematical structure as ‘redundant’ shall follow closely the Principle of Essential and Sufficient Autonomy (PESA) [21], which posits that we should seek the minimal ontology that closes the algebra of dynamical observables. This viewpoint is in-line with the ideas of Leibniz, whose Principle of the Identity of Indiscernibles (PII) motivates a theoretical framework, in which only those elements which affect the dynamics of physically-observable quantities are considered [22, 23].
The second facet of our minimalist approach - the elimination of the structure we have labeled as redundant - will rely upon the identification of ‘dynamical similarities’ [24]; these are a class of non-strictly canonical transformations, which leave the dynamics of the observable degrees of freedom invariant. Theories which possess a dynamical similarity may be subjected to a symmetry reduction procedure, in which a single redundant degree of freedom is excised from the ontology. The main objective of the present work is to generalise this process, which is well-established for classical theories of particles, to a field-theoretic context.
This article is structured in the following way: in section (II), we begin by introducing the necessary mathematical framework of multisymplectic geometry, which we apply to the construction of both the Lagrangian and Hamiltonian formulations of classical field theory. We then present a similar treatment of multicontact geometry; together, these constitute the preliminary geometrical ideas required to precisely formulate both conservative and dissipative field theories.
Following this mathematical introduction, we present the concept of dynamical similarity, as defined for theories of particles, constructed on the (co)tangent bundle of some smooth configuration space manifold. As mentioned above, the identification of a dynamical similarity allows a symmetry reduction procedure to be carried out, the details of which are described at the beginning of section (IV). The remainder of this section is then dedicated to presenting a novel generalisation of the reduction procedure for classical field theories, described within the multisymplectic framework.
Dynamical similarities have been studied in a field-theoretic context in [25]; however, this analysis was conducted within the framework of $k$-symplectic geometry [26], which has a number of disadvantages, as compared to the multisymplectic formalism presented here. Of particular note is that, whilst the vast majority of our results focus on regular systems, the framework we construct is a precursor to a full treatment of contact-reduced gauge theories. The fibred manifolds of multisymplectic geometry are the most natural arena in which to develop such a scheme, and thus justifies our choice of formalism. Furthermore, our principal motivation for developing a rigorous method for treating singular theories lies in the study of General Relativity, where it is known that the Einstein-Hilbert action admits a scaling symmetry; we provide further details of this in section (IV.1). Finally, we close with a number of examples, discussing the relevance and practical implications of our results, before providing several open questions and lines of further investigation.
@@ II.1 Multisymplectic Manifolds and Multivector Fields
If $\Omega$ is closed but 1-degenerate, we shall refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold. Note that a multisymplectic manifold with $k=2$ is simply a standard symplectic manifold. From a field-theoretic point of view, a particularly relevant class of multisymplectic manifold are the first jet bundles over a given ($n$-dimensional differentiable) base manifold, endowed with the Poincaré-Cartan $(n+1)$-forms [^1] The construction and utility of these manifolds shall become clear in due course. [31].
Given a multisymplectic manifold $(\mathcal{M},\Omega)$, sections of the $k^{\textrm{th}}$ exterior power of the tangent bundle $T\mathcal{M}$ define a class of objects known as multivector fields of degree $k$ (or simply $k$-multivector fields) on $\mathcal{M}$. We denote the space of all such multivector fields $\mathfrak{X}^{k}(\mathcal{M}):=\Gamma(\mathcal{M},\wedge^{k}\,T\mathcal{M})$. A $k$-multivector field $\boldsymbol{X}\in\mathfrak{X}^{k}(\mathcal{M})$ is said to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{k}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
@@
We say that an $m$-dimensional distribution $\mathcal{D}\subset T\mathcal{M}$ is locally associated to a non-zero $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$, if there exists some connected open set $\mathcal{V}\subset\mathcal{M}$, such that $\boldsymbol{X}|_{\mathcal{V}}\in\Gamma(\mathcal{V},\wedge^{m}\mathcal{D}|_{\mathcal{V}})$. Note that it is possible for two multivector fields $\boldsymbol{X},\boldsymbol{Y}\in\mathfrak{X}^{m}(\mathcal{M})$ to be locally associated to the same distribution; in such a case, there exists some function $f\in C^{\infty}(\mathcal{V})$, defined on the common connected open subset $\mathcal{V}$, such that $\boldsymbol{Y}=f\boldsymbol{X}$. We shall declare any such multivector fields to be equivalent, and henceforth work with the equivalence classes
$\{\boldsymbol{X}\}_{\mathcal{V}}=\{\boldsymbol{Y}\in\Gamma(\wedge^{m}\mathcal{D}|_{\mathcal{V}})\;|\;\boldsymbol{Y}=f\boldsymbol{X}\;;\;f\in C^{\infty}(\mathcal{V})\,\}$
We refer to a multivector field $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$ as integrable if its locally associated distribution $\mathcal{D}$ is integrable; note that all other multivector fields belonging to the equivalence class of $\boldsymbol{X}$ are automatically integrable, and further, they possess the same integral manifolds [32].
When studying classical field theories, the cases of greatest geometrical interest are those in which we have some bundle structure $\pi:E\rightarrow M$; often this will correspond to the first jet bundle, with $M$ representing the spacetime manifold; however, we shall, for now, retain a somewhat more general setup. We are then interested in the case when the integral manifolds of a given distribution coincide with sections of $\pi$. By this, we mean that, given a section $s:M\rightarrow E$, we shall consider only those integral manifolds $\mathcal{S}\subset E$ which may be expressed as $\mathcal{S}=s(U)$, with $U\subset M$.
Intuitively, this requirement gives rise to a set of integral manifolds which ‘cover’ the total space, in a manner which is transverse to each of the fibres. This notion is made somewhat more precise by introducing the concept of $\pi$-transversality: a multivector field $\boldsymbol{X}\in\mathfrak{X}^{m}(E)$ is said to be $\pi$-transverse if, at each point $p\in E$
$\iota_{\scriptscriptstyle\boldsymbol{X}}(\pi^{*}\xi)=0$
for all $\xi\in\Omega^{m}(M)$ which satisfy $\xi(\pi(p))\neq 0$. Here, the contraction between a locally decomposable $m$-multivector field $\boldsymbol{X}=X_{1}\wedge\,\cdots\,\wedge X_{m}$ and a differential $k$-form $\Xi\in\Omega^{k}(E)$ is naturally defined as
@@
The key statement relating to $\pi$-transverse multivector fields is the following: if $\boldsymbol{X}\in\mathfrak{X}^{m}(E)$ is a locally decomposable integrable multivector field, then it is $\pi$-transverse if and only if its integral manifolds are local sections of $\pi$ [30, 32].
@@ II.2 Lagrangian Field Theory
Having introduced the majority of the mathematical background pertaining to multisymplectic geometry, we shall now illustrate how these ideas are implemented in the Lagrangian description of classical field theories. The fundamental elements of such theories comprise a fibre bundle $\pi:E\rightarrow M$ over the spacetime manifold $M$; we take $M$ to be $m$-dimensional and orientatable, with volume form $\omega$. The total space $E$ is then an $(n+m)$-dimensional manifold, which we refer to as the covariant configuration space; for the vast majority of relevant examples, it will be sufficient to consider total spaces of the form $E=M\times C$, in which $C$ is identified as the $n$-dimensional ‘field space’. Finally, in addition to the bundle $\pi:E\rightarrow M$, we also require the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ [33, 34]; this becomes a manifold over $M$ upon introducing the map $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$. The dimension of $J^{1}E$ is $m+n+nm$, and local coordinates adapted to the bundle structure will be written as $(x^{\mu},y^{a},y^{a}_{\mu})$, with $1\leqslant\mu\leqslant m$ and $1\leqslant a\leqslant n$. Often, when our spacetime manifold is equipped with a Lorentzian metric, we shall adopt the alternative convention that $\textrm{dim}\,M=d$, and $0\leqslant\mu\leqslant d-1$; this will be made clear within the context of the relevant section.
These geometrical elements allow us to define a Lagrangian system on $J^{1}E$, commencing with the introduction of the Lagrangian density $\mathcal{L}$, which is a smooth bundle morphism
$\mathcal{L}:J^{1}E\;\longrightarrow\;\wedge^{m}M$
Since $M$ is orientable, we may consider the Lagrangian density to be a $\widehat{\pi}-$semibasic $m$-form on $J^{1}E$. In general, a differential form $\xi\in\Omega^{m}(J^{1}E)$ is said to be $\widehat{\pi}-$semibasic if $\xi(X_{1},\,\cdots\,,X_{m})=0$ when any one of the vectors $X_{i}\in\mathfrak{X}^{\infty}(J^{1}E)$ belongs to the vertical subspace of $\widehat{\pi}$, that is when $X_{i}\in\textrm{ker}\,\widehat{\pi}_{*}$ [35]. In local bundle coordinates, we shall write
@@
in which $L:J^{1}E\rightarrow\mathbb{R}$ is referred to as the Lagrangian function, and $\widehat{\pi}^{*}\omega$ is the volume form on $M$, pulled back to $J^{1}E$.
In complete analogy to the tangent bundle, upon which we study the standard Lagrangian formulation of particles [36, 37], the space $J^{1}E$ is endowed with a canonical structure: the vertical endomorphism [38, 39, 40]; this is used to define the Cartan forms $\Theta_{L}\in\Omega^{m}(J^{1}E)$ and $\Omega_{L}\in\Omega^{m+1}(J^{1}E)$ [41]. In local bundle coordinates $(x^{\mu},y^{a},y^{a}_{\mu})$, we have
$\Theta_{L}=\frac{\partial L}{\partial y^{a}_{\mu}}\,dy^{a}\wedge d^{m-1}x_{\mu}-\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{m}x\quad\quad\quad\quad\Omega_{L}=-\,d\Theta_{L}$ (2)
The quantity $d^{m-1}x_{\mu}$ is defined in the obvious manner as $d^{m-1}x_{\mu}:=\iota_{\scriptscriptstyle\partial_{\mu}}d^{m}x$. We shall then call the pair $(J^{1}E,\Omega_{L})$ a Lagrangian system, and further specify this to be regular if $\Omega_{L}$ is multisymplectic, and singular if $\Omega_{L}$ is pre-multisymplectic [42].
From the Poincaré-Cartan $m$-form (2), we define the objects $\theta_{L}^{\mu}$ according to
$\theta^{\mu}_{L}=-\,\iota_{\partial_{m}}\,\cdots\;\iota_{\partial_{1}}(\Theta_{L}\wedge dx^{\mu})=\frac{\partial L}{\partial y^{\mu}_{a}}\,dy^{a}$ (3)
Note that these are 1-forms which carry a spacetime index; the $m$-form $\Theta_{L}$ may now be written more succinctly as
$\Theta_{L}=\theta_{L}^{\mu}\wedge d^{m-1}x_{\mu}-E_{L}\,d^{m}x$ (4)
in which $E_{L}\in C^{\infty}(J^{1}E)$ is the energy function corresponding to the Lagrangian $\mathcal{L}$.
The field equations for a Lagrangian system $(J^{1}E,\Omega_{L})$ may be derived via a variational principle (details of which can be found in [31], for example); in the current context however, it suffices to acknowledge that this variational principle seeks a critical section $\phi\in\Gamma(M,E)$ determined by the following criterion
$\phi\;\,\textrm{critical}\quad\quad\Longleftrightarrow\quad\quad\frac{d}{dt}\biggr|_{t=0}\,\int_{M}(j^{1}\phi_{t})\Theta_{L}=0$
Here the variations $\phi_{t}=\sigma_{t}\hskip 0.19919pt\circ\hskip 0.19919pt\phi$ are constructed using a one-parameter family of transformations which generate any compactly-supported vector field in the vertical bundle $\mathfrak{X}^{V(\pi)}(E)$. The section $\phi$ may be represented in local coordinates as $\phi(x)=(x^{\mu},y^{a}(x))$. In the above integral, we are interested in the canonical lifting (also referred to as the first jet prolongation) of $\phi$ to $J^{1}E$, denoted $j^{1}\phi$; this is given locally by
@@
For the purpose of analysing dynamical similarities within a field-theoretic framework, we shall be more interested in a geometrical formulation of the equations of motion, which is entirely equivalent to the above variational approach. In order to present this geometrical formalism, we must expand upon a number of points from section (II.1). Given an $m$-dimensional, $\widehat{\pi}-$transverse, integrable distribution $\mathcal{D}$, we know that it is possible to associate to it an equivalence class of multivector fields $\{\boldsymbol{X}\}$; in local adapted coordinates, a representative of this equivalence class may be expressed as
$\boldsymbol{X}=\bigwedge\limits_{\mu=1}^{m}f\left(\frac{\partial}{\partial x^{\mu}}+F_{\mu}^{a}\,\frac{\partial}{\partial y^{a}}+K_{\mu\nu}^{a}\,\frac{\partial}{\partial y^{a}_{\nu}}\right)$ (6)
in which $f\in C^{\infty}(J^{1}E)$ is some arbitrary (but non-vanishing) function. In almost all practical examples, the condition of $\widehat{\pi}-$transversality is most readily imposed by demanding that $\iota_{\scriptscriptstyle\boldsymbol{X}}(\widehat{\pi}^{*}\omega)=1$, which sets $f=1$ in the above expression.
Suppose that $\psi\in\Gamma(M,J^{1}E)$ is a section of $\widehat{\pi}$, which we choose to represent locally as $\psi(x)=(x^{\mu},y^{a}(x),y^{a}_{\mu}(x))$; we shall refer to this as an integral section of the multivector field $\boldsymbol{X}$ if $F_{\mu}^{a}=y^{a}_{\mu}$ in the expansion (6). Additionally, an integral section is said to be holonomic if it is the canonical lifting of a section $\phi\in\Gamma(M,E)$, that is to say, if we may write $\psi=j^{1}\phi$. A locally decomposable multivector field is then called holonomic if its integral sections are all holonomic. If a locally decomposable multivector field $\boldsymbol{X}$ has the form (6), but is not integrable, then we say that it is semi-holonomic.
The above discussion is quite general; let us now see how these ideas apply in the context of the Lagrangian field equations. Suppose that $\phi\in\Gamma(M,E)$ is a critical section for the Lagrangian system $(J^{1}E,\Omega_{L})$, as determined by the variational principle. Then the canonical lifting $j^{1}\phi$ satisfies
$(j^{1}\phi)^{*}\,\iota_{\scriptscriptstyle X}\Omega_{L}=0\quad\quad\quad\textrm{for all}\;X\in\mathfrak{X}^{\infty}(J^{1}E)$ (7)
Additionally, the section $j^{1}\phi$, expressed in local coordinates as $j^{1}\phi(x)=\left(x^{\mu},y^{a}(x),\frac{\partial y^{a}}{\partial x^{\mu}}|_{x}\right)$, satisfies the following Euler-Lagrange field equations
Finally, we find that $j^{1}\phi$ is an integral section of a class of holonomic multivector fields $\{\boldsymbol{X}_{L}\}$, all of which satisfy
$\iota_{\scriptscriptstyle\boldsymbol{X}_{L}}\Omega_{L}=0$ (9)
Note that equations (8) and (9) are highly reminiscent of their standard symplectic analogues [43, 44]. Indeed, this must be the case, given that, as mentioned in section (II.1), symplectic manifolds are but a special case of multisymplectic manifolds, for which the 1-non-degenerate form $\Omega$ coincides with the symplectic 2-form [45].

### v4

@@ Abstract
Symmetry under a particular class of non-strictly canonical transformation may be used to identify, and subsequently excise degrees of freedom which do not contribute to the closure of the algebra of dynamical observables. Such redundant degrees of freedom may physically be identified with empirically-inaccessible measures of global scale. In this article, we present a mathematical framework which extends the symmetry reduction procedure to theories of classical fields, in both the Lagrangian and Hamiltonian settings. In order to maintain Lorentz covariance, while simultaneously working with a finite-dimensional phase space, we employ the De-Donder Weyl formalism, for which the natural description is formulated in terms of the fibered manifolds of multisymplectic geometry. We subsequently analyse a number of simple examples, and provide a discussion of the broader implications of our construction.
@@ I Introduction
Our most precise and well-tested description of nature is currently formulated in the language of field theory, both at a classical and quantum level [1]. Indeed, the advent of a field-theoretic viewpoint has arguably been one of the most significant shifts in perspective of the last few centuries. The particular mathematical framework introduced to model the phenomena of interest depends upon context, and, to a lesser extent, convenience and personal preference. From the perspective of the symplectic description of classical field theories, we highlight two philosophically-distinct approaches. The canonical formalism, often employed to construct the quantum theory from the underlying classical framework, studies dynamical evolution, as defined by the initial-value problem. In particular, one sacrifices Lorentz covariance, selecting a privileged time coordinate, and considers field variables defined on spatial slices of fixed temporal parameter [2].
The spacetime-covariant formalism, by contrast, treats events at a single spacetime point, considering all coordinates (including time) equally [3, 4]. Multisymplectic geometry provides a rigorous framework within which to study classical field theories in just such a covariant setting, which we shall refer to as the De-Donder Weyl formalism [5, 6, 7, 8, 9]. While the canonical approach, or alternative tools, such as variational bicomplexes, are often favoured in the literature [10], their infinite-dimensionality presents a significant hindrance to the study of the class of symmetries of interest in the present work [11], which we shall refer to as scaling symmetries or dynamical similarities.
Scaling symmetries arise when our mathematical formalism contains more structure than that required to describe the dynamical evolution of the observable degrees of freedom [12]. Not only is such a scenario undesirable from the perspective of elegance and simplicity, but the superfluous structure can, at times, introduce pathological behaviour that is an artifact of our choice of description, and not reflective of the true nature of the underlying physical system. For example, it is the universally-accepted practice to describe the dynamical evolution of homogeneous, isotropic cosmologies using a dimensionless scale factor $a(t)$ [13]. It is known that the value of the scale factor by itself cannot be determined by experiment; it is only the ratio of $a(t)$ at two different times that encodes physical information. It has been demonstrated that one can formulate an equivalent cosmological description, without ever introducing a scale factor [14, 15]. Moreover, as alluded to above, this description is perfectly well-defined at those points where the singular nature of the scale factor causes the conventional theory to break down [16].
We argue that, following the Principle of Essential and Sufficient Autonomy, when seeking a theoretical description of the natural world, one should first construct a framework which amply describes the phenomena of interest, before progressively reducing its superfluous structure, until a minimally-sufficient yet consistent theory remains [17, 18, 19, 20]. By considering the multisymplectic formulation of classical field theory, introduced in section (II), we demonstrate that invariance under changes to a system’s global scale indicates the presence of a single redundant degree of freedom, corresponding to the generator of these rescalings. The process by which such structure is excised from our theory is known as ‘contact reduction’, and is introduced in section (IV). We then dedicate sections (V) and (VI) to the field-theoretic generalisation of this procedure. Throughout, the multisymplectic framework is indispensable for our construction; it is not clear, for instance, that the more conventional methods admit a means to assign a well-defined structure to our theory, after having eliminated a scaling degree of freedom. By contrast, when the fibered manifolds of the multisymplectic formalism are employed, we shall demonstrate that the reduced theory is defined on a space naturally identified as a multicontact manifold [21, 22, 23].
The ideas presented throughout constitute a necessary preliminary framework that is precursory to a full treatment of contact-reduced gauge theories. Our principal objective in developing methods for treating singular theories lies in the study of General Relativity, where it is known that the Einstein-Hilbert action admits a scaling symmetry [24]. A geometrical constraint algorithm has been developed for singular field theories described using multisymplectic geometry [25, 26], and the interaction between phase space restriction, and the contact reduction procedure presented here should be examined, prior to analysing scaling symmetries of gravitational actions.
Finally, we close with a number of examples, in which a scaling symmetry is identified, and the reduction procedure carried out. These examples have been chosen to be sufficiently involved so as to admit a detailed discussion of the physical implications of our construction. They are, however, simple enough so as to allow readers unfamiliar with multisymplectic geometry to follow, without additional contextual complications. Having presented these examples, we shall indicate a number of more physically-motivated theories to which our formalism is applicable, as well as discussing several open questions and lines of further investigation.
@@ II.1 Multisymplectic Manifolds and Multivector Fields
If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold. Given a multisymplectic manifold $(\mathcal{M},\Omega)$, sections of the $k^{\textrm{th}}$ exterior power of $T\mathcal{M}$ define a class of objects known as multivector fields of degree $k$ on $\mathcal{M}$. We denote the space of all such multivector fields $\mathfrak{X}^{k}(\mathcal{M}):=\Gamma(\mathcal{M},\wedge^{k}\,T\mathcal{M})$. A $k$-multivector field $\boldsymbol{X}\in\mathfrak{X}^{k}(\mathcal{M})$ is said to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{k}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
@@
We say that an $m$-dimensional distribution $\mathcal{D}\subset T\mathcal{M}$ is locally associated to a non-zero $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$, if there exists some connected open set $\mathcal{V}\subset\mathcal{M}$, such that $\boldsymbol{X}|_{\mathcal{V}}\in\Gamma(\mathcal{V},\wedge^{m}\mathcal{D}|_{\mathcal{V}})$, and $\boldsymbol{X}\in\mathfrak{X}^{m}(\mathcal{M})$ is said to be integrable if its locally associated distribution is integrable.
Finally, we define the contraction between a locally decomposable $m$-multivector field $\boldsymbol{X}=X_{1}\wedge\,\cdots\,\wedge X_{m}$ and a differential $k$-form $\Xi\in\Omega^{k}(\mathcal{M})$ as
@@ II.2 Lagrangian Field Theory
We begin by introducing a fibre bundle $\pi:E\rightarrow M$ over the orientable, $d$-dimensional spacetime manifold $M$, with volume form $\omega$, and local coordinates $x^{\mu}$ ($\mu=0,\,\cdots,d-1$). The total space $E$ is a $(d+n)$-dimensional manifold, referred to as the covariant configuration space, upon which local coordinates are written $(x^{\mu},y^{a})$, with $1\leqslant a\leqslant n$. Introducing the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ [31, 32], together with the map $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$, the Lagrangian density is a $\widehat{\pi}-$semibasic $d$-form on $J^{1}E$, expressed as
@@
in which $L:J^{1}E\rightarrow\mathbb{R}$ denotes the Lagrangian function. In local coordinates $(x^{\mu},y^{a},y^{a}_{\mu})$ on $J^{1}E$, the volume form simply reads $\widehat{\pi}^{*}\omega=d^{d}x$. From the Lagrangian function, we define two Cartan forms $\Theta_{L}\in\Omega^{d}(J^{1}E)$ and $\Omega_{L}\in\Omega^{d+1}(J^{1}E)$ as
$\Theta_{L}=\frac{\partial L}{\partial y^{a}_{\mu}}\,dy^{a}\wedge d^{d-1}x_{\mu}-\left(\frac{\partial L}{\partial y^{a}_{\mu}}\,y^{a}_{\mu}-L\right)\,d^{d}x\quad\quad\quad\quad\Omega_{L}:=-\,d\Theta_{L}$ (2)
The quantity $d^{d-1}x_{\mu}$ is defined in the obvious manner as $d^{d-1}x_{\mu}:=\iota_{\scriptscriptstyle\partial_{\mu}}d^{d}x$. We refer to the pair $(J^{1}E,\Omega_{L})$ as a Lagrangian system, further specifying this to be regular if $\Omega_{L}$ is multisymplectic, and singular if it is pre-multisymplectic [26]. From the Cartan $d$-form $\Theta_{L}$, we define the following objects, used extensively in our analysis of scaling symmetries
$\theta^{\mu}_{L}:=-\,\iota_{\partial_{d-1}}\,\cdots\;\iota_{\partial_{0}}(\Theta_{L}\wedge dx^{\mu})=\frac{\partial L}{\partial y^{\mu}_{a}}\,dy^{a}$ (3)
The field equations for a Lagrangian system $(J^{1}E,\Omega_{L})$ may be derived from a variational principle (details of which can be found in [33]); here, it suffices to note that this variational principle identifies critical sections $\phi\in\Gamma(M,E)$, whose canonical lifting $j^{1}\phi$ to $J^{1}E$ satisfy the following Euler-Lagrange field equations
If $\phi$ is expressed locally as $\phi(x)=\left(x^{\mu},y^{a}(x)\right)$, then
@@
In general, a section $\psi:M\rightarrow J^{1}E$ is said to be holonomic if it may be expressed as the canonical lifting $\psi=j^{1}\chi$ of some $\chi:M\rightarrow E$. An integrable multivector field whose integral sections all have this property is itself referred to as holonomic. The field equations may be formulated geometrically seeking locally decomposable, $\widehat{\pi}-$transverse multivector fields $\boldsymbol{X}_{L}$ which satisfy $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$. Such multivector fields are expressed locally as
$\boldsymbol{X}_{L}=\bigwedge\limits_{\mu=0}^{d-1}f\left(\frac{\partial}{\partial x^{\mu}}+F_{\mu}^{a}\,\frac{\partial}{\partial y^{a}}+K_{\mu\nu}^{a}\,\frac{\partial}{\partial y^{a}_{\nu}}\right)$ (6)
for some non-vanishing $f\in C^{\infty}(J^{1}E)$. In practice, $\widehat{\pi}-$transversality is most easily enforced by demanding that $\iota_{\scriptscriptstyle\boldsymbol{X}}(\widehat{\pi}^{*}\omega)=1$, which sets $f$ to unity. When $\boldsymbol{X}_{L}$ is integrable, it is holonomic only when $F_{\mu}^{a}=y^{a}_{\mu}$ in the above expansion; if $F_{\mu}^{a}=y^{a}_{\mu}$, but $\boldsymbol{X}_{L}$ is not integrable, then we say that it is semi-holonomic.